Good day!Please help me to answer this.I was told by my teacher to give an equation to this:
Five years ago,John's age was half of the age he will be in 8 years.How old is he now?
Thank you! May God bless you!

Answer by stanbon(75887)

Five years ago,John's age was half of the age he will be in 8 years.How old is he now?
------------------
Let John's age NOW be "x" yrs old.
----------------
Five years ago he was "x-5" yrs old
-------------------------
In 8 years from now he will be "x+8".
------------------------
EQUATION:
x-5 = (1/2)(x+8)
Multiply both sides by 2 to get:
2x-10 = x+8
x = 18 (John's age NOW is 18.)
==============================
Cheers,
Stan H.